OpenStudy (anonymous):

Price of the item: $12.99 what is the item price with a 5 % sales tax.

OpenStudy (anonymous):

\(12.99\times 1.05\) and a calculator

OpenStudy (anonymous):

do i use part=percent x whole

Parth (parthkohli):

Calculator's for newbs.

Parth (parthkohli):

I approximate stuff.

OpenStudy (anonymous):

you can estimate 10% of 13 is 1.3 so 5% of 13 is .65

Parth (parthkohli):

\[12.99(1 + 0.05) = (13 - 0.01)(1 + 0.05)\]FOIL.

Parth (parthkohli):

That one is, if you don't want to approximate. ^

OpenStudy (anonymous):

i think approximating is easier

Parth (parthkohli):

It always is.

OpenStudy (anonymous):

soo do i add .65 to 12.99

Parth (parthkohli):

Yes.